# On Solving Asymmetric Diagonally Dominant Linear Systems in Sublinear Time

### v1

@@ 7 The Bidirectional Method
So, instead of directly using $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ as as estimate of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$, we estimate the right-hand side of the above equation using random-walk samplings from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$ to reduce approximation error. We employ the same sampling scheme as in Algorithm 3 to obtain a walk length $\ell$ and the coordinate $v$ reached by the random walk after $\ell$ steps, but each sampled value now involves the summation $\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$. We take the average of the estimates across $n_{\mathrm{s}}$ independent samples and add it to $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ to obtain the final estimate $\hat{x}$.
We note that, compared to the sampling scheme in [CWW25], our approach eliminates the need for additional data structures to maintain the prefix sums of residues, since we can directly compute $\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$ in $O(L)$ time per sample without increasing the asymptotic complexity.
@@ 12    $v\leftarrow v^{\prime}$
The sum of $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ and each sampled value in the bidirectional method gives an unbiased estimate of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$.
@@
We use Algorithm 5. Note that each sampled value equals $\sigma\cdot\frac{1}{2}\|\boldsymbol{t}\|_{1}L\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$ for some $\sigma\in\{0,\pm 1\}$ and $v\in[n]$ and the Push algorithm ensures that $\left\|\boldsymbol{r}^{(\ell^{\prime})}\right\|_{\infty}\leq r_{\max}$ for each $\ell^{\prime}\in[0,L-2]$. Thus, the absolute value of each sampled value is at most $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}$. Using the Hoeffding bound, it follows that
@@ 8.1 Understanding ForwardPush and BackwardPush
The result follows by verifying that the described variable transformations convert the procedure of $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\boldsymbol{e}_{s},L,r_{\max}\right)$ into the procedure of $\textup{{ForwardPush}}(G,s,\alpha,L,r_{\max})$ exactly. For the initialization, Push sets $\boldsymbol{r}^{(0)}(s)=\alpha/d^{+}_{G}(s)$, so $\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)=\alpha$, ,matching the initialization in ForwardPush. The condition for performing the push operation on a coordinate $v$ at level $\ell$ in Push is $\left|\boldsymbol{r}^{(\ell)}(v)\right|>r_{\max}$, which is equivalent to $\frac{\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)}{d^{+}_{G}(v)}>r_{\max}$ since the residues in ForwardPush are nonnegative, matching the condition in ForwardPush. In the push operation on $v$ at level $\ell$, Push increments $\boldsymbol{r}^{(\ell+1)}(u)$ by $\frac{(1-\alpha)\mathbf{A}_{G}(v,u)}{2d^{+}_{G}(u)}\cdot\boldsymbol{r}^{(\ell)}(v)$ for each $u$ with $\mathbf{A}_{G}(v,u)>0$. This is equivalent to incrementing $\boldsymbol{r}_{\mathrm{FP}}^{(\ell+1)}(u)$ by $d^{+}_{G}(u)\cdot\frac{(1-\alpha)\mathbf{A}_{G}(v,u)}{2d^{+}_{G}(u)}\cdot\boldsymbol{r}^{(\ell)}(v)=\frac{1}{2}(1-\alpha)\mathbf{A}_{G}(v,u)\cdot\boldsymbol{r}^{(\ell)}(v)=\frac{1}{2}(1-\alpha)\cdot\frac{\mathbf{A}_{G}(v,u)}{d^{+}_{G}(v)}\cdot\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)$ for each out-neighbor $u$ of $v$, matching the corresponding step in ForwardPush. This verifies the equivalence of the two algorithms. ∎
@@
When $G$ is Eulerian, the outdegree matrix of $G^{\top}$ equals $\mathbf{D}_{G}$. Thus, by the previous claims, ForwardPush on $G$ corresponds to $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\boldsymbol{e}_{s},L,r_{\max}\right)$, while BackwardPush on $G^{\top}$ with threshold $r_{\max}\cdot d_{G}(s)$ is equivalent to $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\mathbf{D}_{G}\boldsymbol{e}_{s},L,r_{\max}\cdot d_{G}(s)\right)$. In the two sets of parameters for Push, changing $\boldsymbol{b}=\alpha\boldsymbol{e}_{s}$ to $\boldsymbol{b}=\alpha\mathbf{D}_{G}\boldsymbol{e}_{s}$ scales the initial residue $\boldsymbol{r}^{(0)}(s)$ by $d_{G}(s)$, which in term scales all the reserves and residues by $d_{G}(s)$. As the threshold $r_{\max}$ is also scaled by $d_{G}(s)$, the procedures of the two algorithms remain equivalent. ∎
@@ 8.2 Results for PageRank Computation when $\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top}$ is RCDD
To prove Theorem 1.9, we also establish some lower bounds on PageRank values in the next lemma, which may be of independent interest. This lemma is partly inspired by [BPP23, Lemma 5.13], [WW23, Wan24], and [WWWY24, Theorem 1.1].
@@
Note that this complexity can also be bounded by $\widetilde{O}\left(1/\varepsilon^{2}\right)$ since $\boldsymbol{\pi}_{G,\alpha}(t)\geq\alpha/n$. This matches a result given in a concurrent work [TWWY25], which states that if $(1-\alpha)\Delta^{-}_{G}\leq 1$ for an unweighted graph $G$, then we can approximate single-node PageRank within relative error $\varepsilon$ in $\widetilde{O}\left(1\right)$ time, for constant $\alpha$ and $\varepsilon$. Compared to their algorithm, our algorithm is simpler and only performs random-walk sampling from $t$.
@@ 8.3 A Lower Bound on the Accuracy Parameter for SDD Solvers
Suppose that there exists a randomized algorithm that computes an estimate $\hat{x}_{t}$ such that $\Pr\left\{\big{|}\hat{x}_{t}-\boldsymbol{x}^{\ast}(t)\big{|}\leq\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}\right\}\geq\frac{3}{4}$ for any SDD system $\mathbf{S}\boldsymbol{x}=\boldsymbol{b}$ in $O\left(\gamma^{-\nu}\varepsilon^{-\tau}\right)$ time. Then there exists a randomized algorithm that, given $\delta_{G}$, estimates $\boldsymbol{\pi}_{G,\alpha}(t)$ on undirected unweighted graphs $G$ within constant relative error and with success probability at least $3/4$ in time $O\left(\big{(}d_{G}(t)/\delta_{G}\big{)}^{\tau}/\alpha^{\nu+\tau}\right)$.
@@
[Wan24] establishes a complexity lower bound of $\Omega\big{(}d_{G}(t)/\delta_{G}\big{)}$ for estimating $\boldsymbol{\pi}_{G,\alpha}(t)$ within constant relative error with constant success probability on undirected unweighted graphs, where $\alpha$ is constant and the bound holds for any possible combination of $\delta_{G}$ and $d_{G}(t)$. This lower bound applies to the number of queries to the graph structure. Therefore, combining this lower bound with Lemma 8.7 and noting that the reduction uses $\varepsilon:=\Theta\big{(}\alpha\delta_{G}/d_{G}(t)\big{)}=\Omega(1/n)$ yield the desired lower bound of $\Omega(1/\varepsilon)$ for $\varepsilon=\Omega(1/n)$. ∎
@@ 9 Connections with Effective Resistance Computation
Now we can directly apply Theorems 1.4, 1.7, and 1.8 to prove Corollary 1.10. The only remaining detail in the proof is to derive a better setting of $L$ for the case $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$.
@@
Following the proof of Theorem 1.3, we can derive that $\left|\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}\right|\leq\frac{1}{2\gamma}\cdot e^{-\gamma L}\cdot\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{t}\right\|_{2}\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{b}\right\|_{2}$. As $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$, we have $\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{t}\right\|_{2}\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{b}\right\|_{2}=\frac{1}{d_{G}(s)}+\frac{1}{d_{G}(t)}$. Thus, setting $L:=\Theta\left(\frac{1}{\gamma}\log\left(\frac{1}{\gamma\varepsilon}\left(\frac{1}{d_{G}(s)}+\frac{1}{d_{G}(t)}\right)\right)\right)$ ensures that $\left|\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}\right|\leq\frac{1}{2}\varepsilon$. The corollary then follows by applying Theorems 1.4, 1.7, and 1.8 with this setting of $L$ and noting that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}=1/\min\big{(}d_{G}(s),d_{G}(t)\big{)}$ and $f_{\mathrm{row}}(\mathbf{M})=O(1)$ in this case. ∎
@@ 10 Acknowledgments
This research was supported by National Natural Science Foundation of China (No. 92470128, No. U2241212). Mingji Yang wishes to thank Guanyu Cui for helpful discussions on effective resistance.

### v2

@@ 7 The Bidirectional Method
So, instead of directly using $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ as an estimate of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$, we estimate the right-hand side of the above equation using random-walk samplings from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$ to reduce approximation error. We employ the same sampling scheme as in Algorithm 3 to obtain a walk length $\ell$ and the coordinate $v$ reached by the random walk after $\ell$ steps, but each sampled value now involves the summation $\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$. We take the average of the estimates across $n_{\mathrm{s}}$ independent samples and add it to $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ to obtain the final estimate $\hat{x}$.
We note that, compared to the sampling scheme in [CWW25], our approach does not need additional data structures to maintain the prefix sums of residues, since we can directly compute $\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$ in $O(L)$ time per sample without increasing the asymptotic complexity.
@@ 12    $v\leftarrow v^{\prime}$
The sum of $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ and each sampled value in the bidirectional method described above gives an unbiased estimate of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$.
@@
We use the bidirectional method as in Algorithm 5. Note that each sampled value equals $\sigma\cdot\frac{1}{2}\|\boldsymbol{t}\|_{1}L\sum_{\ell^{\prime}=0}^{\min(L-\ell-1,L-2)}\boldsymbol{r}^{(\ell^{\prime})}(v)$ for some $\sigma\in\{0,\pm 1\}$ and $v\in[n]$ and the Push algorithm ensures that $\left\|\boldsymbol{r}^{(\ell^{\prime})}\right\|_{\infty}\leq r_{\max}$ for each $\ell^{\prime}\in[0,L-2]$. Thus, the absolute value of each sampled value is at most $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}$. Using the Hoeffding bound, it follows that
@@ 8.1 Understanding ForwardPush and BackwardPush
The result follows by verifying that the described variable transformations convert the procedure of $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\boldsymbol{e}_{s},L,r_{\max}\right)$ into the procedure of $\textup{{ForwardPush}}(G,s,\alpha,L,r_{\max})$ exactly. For the initialization, Push sets $\boldsymbol{r}^{(0)}(s)=\alpha/d^{+}_{G}(s)$, so $\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)=\alpha$, matching the initialization in ForwardPush. The condition for performing the push operation on a coordinate $v$ at level $\ell$ in Push is $\left|\boldsymbol{r}^{(\ell)}(v)\right|>r_{\max}$, which is equivalent to $\frac{\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)}{d^{+}_{G}(v)}>r_{\max}$ since the residues in ForwardPush are nonnegative, matching the condition in ForwardPush. In the push operation on $v$ at level $\ell$, Push increments $\boldsymbol{r}^{(\ell+1)}(u)$ by $\frac{(1-\alpha)\mathbf{A}_{G}(v,u)}{2d^{+}_{G}(u)}\cdot\boldsymbol{r}^{(\ell)}(v)$ for each $u$ with $\mathbf{A}_{G}(v,u)>0$. This is equivalent to incrementing $\boldsymbol{r}_{\mathrm{FP}}^{(\ell+1)}(u)$ by $d^{+}_{G}(u)\cdot\frac{(1-\alpha)\mathbf{A}_{G}(v,u)}{2d^{+}_{G}(u)}\cdot\boldsymbol{r}^{(\ell)}(v)=\frac{1}{2}(1-\alpha)\mathbf{A}_{G}(v,u)\cdot\boldsymbol{r}^{(\ell)}(v)=\frac{1}{2}(1-\alpha)\cdot\frac{\mathbf{A}_{G}(v,u)}{d^{+}_{G}(v)}\cdot\boldsymbol{r}_{\mathrm{FP}}^{(\ell)}(v)$ for each out-neighbor $u$ of $v$, matching the corresponding step in ForwardPush. This verifies the equivalence of the two algorithms. ∎
@@
When $G$ is Eulerian, the outdegree matrix of $G^{\top}$ equals $\mathbf{D}_{G}$. Thus, by the previous claims, ForwardPush on $G$ corresponds to $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\boldsymbol{e}_{s},L,r_{\max}\right)$, while BackwardPush on $G^{\top}$ with threshold $r_{\max}\cdot d_{G}(s)$ is equivalent to $\textup{{Push}}\left(\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top},\alpha\mathbf{D}_{G}\boldsymbol{e}_{s},L,r_{\max}\cdot d_{G}(s)\right)$. In the two sets of parameters for Push, changing $\boldsymbol{b}=\alpha\boldsymbol{e}_{s}$ to $\boldsymbol{b}=\alpha\mathbf{D}_{G}\boldsymbol{e}_{s}$ scales the initial residue $\boldsymbol{r}^{(0)}(s)$ by $d_{G}(s)$, which in turn scales all the reserves and residues by $d_{G}(s)$. As the threshold $r_{\max}$ is also scaled by $d_{G}(s)$, the procedures of the two algorithms remain equivalent. ∎
@@ 8.2 Results for PageRank Computation when $\mathbf{D}_{G}-(1-\alpha)\mathbf{A}_{G}^{\top}$ is RCDD
To prove Theorem 1.9, we establish some lower bounds on PageRank values in the next lemma, which may be of independent interest. This lemma is partly inspired by [BPP23, Lemma 5.13], [WW23, Wan24], and [WWWY24, Theorem 1.1].
@@
Note that this complexity can also be bounded by $\widetilde{O}\left(1/\varepsilon^{2}\right)$ since $\boldsymbol{\pi}_{G,\alpha}(t)\geq\alpha/n$. This matches a result given in a concurrent work [TWWY26], which states that if $(1-\alpha)\Delta^{-}_{G}\leq 1$ for an unweighted graph $G$, then we can approximate single-node PageRank within relative error $\varepsilon$ in $\widetilde{O}\left(1\right)$ time, for constant $\alpha$ and $\varepsilon$. Compared to their algorithm, our algorithm is simpler and only performs random-walk sampling from $t$.
@@ 8.3 A Lower Bound on the Accuracy Parameter for SDD Solvers
Suppose that there exists a randomized algorithm that computes an estimate $\hat{x}_{t}$ such that $\Pr\left\{\big|\hat{x}_{t}-\boldsymbol{x}^{\ast}(t)\big|\leq\varepsilon\|\boldsymbol{x}^{\ast}\|_{\infty}\right\}\geq\frac{3}{4}$ for any SDD system $\mathbf{S}\boldsymbol{x}=\boldsymbol{b}$ in $O\left(\gamma^{-\nu}\varepsilon^{-\tau}\right)$ time. Then there exists a randomized algorithm that, given $\delta_{G}$, estimates $\boldsymbol{\pi}_{G,\alpha}(t)$ on unweighted undirected graphs $G$ within constant relative error and with success probability at least $3/4$ in time $O\left(\big(d_{G}(t)/\delta_{G}\big)^{\tau}/\alpha^{\nu+\tau}\right)$.
@@
[Wan24] establishes a complexity lower bound of $\Omega\big(d_{G}(t)/\delta_{G}\big)$ for estimating $\boldsymbol{\pi}_{G,\alpha}(t)$ within constant relative error with constant success probability on unweighted undirected graphs, where $\alpha$ is constant and the bound holds for any possible combination of $\delta_{G}$ and $d_{G}(t)$. This lower bound applies to the number of queries to the graph structure. Therefore, combining this lower bound with Lemma 8.7 and noting that the reduction uses $\varepsilon:=\Theta\big(\alpha\delta_{G}/d_{G}(t)\big)=\Omega(1/n)$ yield the desired lower bound of $\Omega(1/\varepsilon)$ for $\varepsilon=\Omega(1/n)$. ∎
@@ 9 Connections with Effective Resistance Computation
Now we can directly apply Theorems 1.4 and 1.8 to prove Corollary 1.10. The only remaining detail in the proof is to derive a better setting of $L$ for the case $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$.
@@
Following the proof of Theorem 1.3, we can derive that $\left|\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}\right|\leq\frac{1}{2\gamma}\cdot e^{-\gamma L}\cdot\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{t}\right\|_{2}\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{b}\right\|_{2}$. As $\mathbf{M}=\mathbf{L}_{G}$ and $\boldsymbol{b}=\boldsymbol{t}=\boldsymbol{e}_{s}-\boldsymbol{e}_{t}$, we have $\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{t}\right\|_{2}\left\|\mathbf{D}_{\mathbf{M}}^{-1/2}\boldsymbol{b}\right\|_{2}=\frac{1}{d_{G}(s)}+\frac{1}{d_{G}(t)}$. Thus, setting $L:=\Theta\left(\frac{1}{\gamma}\log\left(\frac{1}{\gamma\varepsilon}\left(\frac{1}{d_{G}(s)}+\frac{1}{d_{G}(t)}\right)\right)\right)$ ensures that $\left|\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}\right|\leq\frac{1}{2}\varepsilon$. The corollary then follows by applying Theorems 1.4 and 1.8 with this setting of $L$ and noting that $\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}=1/\min\big(d_{G}(s),d_{G}(t)\big)$ and $f_{\mathrm{row}}(\mathbf{M})=O(1)$ in this case. ∎
@@ 10 Acknowledgments
This research was supported by National Natural Science Foundation of China (No. 92470128, No. U2241212). We thank the anonymous reviewers for their valuable comments. Mingji Yang thanks Prof. Shang-Hua Teng for inspiring discussions and encouragement, and Guanyu Cui for helpful discussions on effective resistance.
@@ A.2 Deferred Proofs
range
pages16 rangepages23 rangepages12 rangepages30 rangepages12 rangepages-1 rangepages9 rangepages6 rangepages11 rangepages10 rangepages41 rangepages-1 rangepages12 rangepages10 rangepages10 rangepages12 rangepages-1 rangepages13 rangepages3 rangepages26 rangepages21 rangepages3 rangepages43 rangepages42 rangepages8 rangepages10 rangepages1 rangepages20 rangepages44 rangepages19 rangepages13 rangepages10 rangepages10 rangepages46 rangepages1 rangepages18 rangepages10 rangepages16 rangepages5 rangepages14 rangepages10 rangepages51 rangepages10 rangepages13 rangepages21 rangepages141 rangepages9 rangepages4 rangepages13 rangepages12 rangepages-1 rangepages29 rangepages-1 rangepages21